# Q-BAR: Blogger Anomaly Recognition via Quantum-enhanced Manifold Learning

### v1

@@ 1 Introduction
The digital information ecosystem is currently witnessing a crisis of authenticity that transcends pixel-level forgery (farid2022creating; wardle2017information). While deepfake detection focuses on AI-synthesized artifacts, a more pervasive threat has emerged in the form of semantic mutation (guarnera2020fighting). In this phenomenon, unscrupulous actors, often operating as content farms, use traditional editing tools to reorder, decontextualize, or selectively splice genuine footage (mubarak2023survey). A video subjected to these attacks remains visually authentic, containing the real speaker and genuine audio, but the underlying semantics and intent are fundamentally mutated to fabricate false narratives or clickbait (kishwar2025regulating; luo2021newsclippings).
The societal impact of such manipulation is profound and immediate  (vaccari2020deepfakes). A prominent example occurred on short-video platforms involving the top-tier technology creator MediaStorm (Yingshijufeng) [^1] See, for example, the widely discussed incident regarding unauthorized re-editing and platform content degradation reported in https://www.zhihu.com/question/1972852612236850378 (2025) or MediaStorm’s official response on Bilibili.. Malicious re-editing of their reviews resulted in a complete semantic inversion of the original critique, fabricating conflicts that did not exist. This incident not only dominated platform trending lists but also instigated widespread misinformation and targeted cyber-harassment against the creator. Such events underscore a critical vulnerability in the current media landscape: when visual fidelity is weaponized to distort factual truth, it can rapidly incite irrational public sentiment and cause irreversible reputational damage before the original context can be restored (vosoughi2018spread; vaccari2020deepfakes). Detecting these high-level logic violations requires modeling the unique semantic manifold of the original creator (ruff2018deep). A genuine blogger maintains a consistent latent signature defined by their linguistic patterns, logical structures, and prosodic delivery (stamatatos2009survey). Ideally, an anomaly detector would learn the topology of this manifold and flag any video that drifts into low-density regions, indicating a deviation from the creator’s established persona.
@@
However, applying classical deep learning to this task faces a critical bottleneck: data scarcity (huang2022quantum; huang2021power; wang2025parameter). A specific creator typically possesses a limited repository of high-quality, long-form original videos, often ranging from 20 to 50 samples. Classical anomaly detection methods, such as Deep Support Vector Data Description (Deep SVDD) (ruff2018deep) or Variational Autoencoders (VAE) (park2023variational; an2015variational), are notoriously data-hungry. When trained on such small, high-dimensional multimodal datasets, these models tend to suffer from severe overfitting or require massive parameter sets that fail to generalize to unseen semantic attacks (wang2023quantum). To surmount the obstacle of data scarcity, we turn to the emerging paradigm of Quantum Machine Learning (QML) (biamonte2017quantum). Recent theoretical breakthroughs have rigorously established that quantum models can extract distinct information from complex, high-dimensional feature spaces more efficiently than their classical counterparts (wang2025quantum). Notably, Huang et al. demonstrated that quantum learning algorithms can achieve a substantial advantage in sample complexity, enabling accurate classification with significantly fewer training examples (huang2021power; huang2022quantum). This characteristic renders QML uniquely suitable for modeling the semantic manifolds of individual creators, where the “ground truth” is sparse yet highly complex.
Motivated by these findings, our approach integrates the Parameter-Efficient Quantum Anomaly Detection (PEQAD) algorithm into the domain of multimedia forensics (wang2025parameter). Specifically, we fuse textual, audio, and visual features into a compact representation and employ a variational quantum circuit (VQC) to map normal blogger content into a minimal hypersphere in the quantum state space. This strategic integration yields a threefold enhancement. First, it mitigates overfitting by achieving robust anomaly detection using a few trainable quantum parameters, whereas comparable classical models often require hundreds. Second, the quantum state representation naturally models the entangled relationships between a blogger’s tone and logic (gao2022enhancing). Third, the reduced model complexity facilitates potential deployment on near-term quantum devices or efficient simulators.
@@ 2 Related Work
Traditional VAD identifies abnormal events in surveillance feeds, typically employing autoencoders (hasan2016learning) or multiple instance learning to detect deviations from normal motion patterns (sultani2018real). However, these methods rely heavily on visual cues (optical flow, gradients) (liu2018future) and are agnostic to linguistic semantics. They do not focus on detecting a video where a person is speaking calmly but saying something successfully spliced to reverse their original meaning (abdalla2025video).
@@
Research in misinformation has largely focused on “Fake News” classification via textual features or graph-based propagation patterns. Meanwhile, Deepfake detection analyzes frequency domain artefacts or blending boundaries in synthesized faces (rossler2019faceforensics). Our work addresses a distinct category: Cheapfakes or shallowfakes, where the manipulation is editorial rather than generative. As noted by (paris2019deepfakes), this semantic manipulation is often more dangerous and harder to detect than AI-generated content. Crucially, existing deepfake detectors fail on these semantic alterations because no pixel-level artefacts are present.
@@
Recent works in NLP have explored logical consistency in text summarization (maynez2020faithfulness). We extend this notion to the multimodal domain, treating the blogger’s historical corpus as the “ground truth” reference for consistency checking.
Quantum Machine Learning (QML).
Quantum Machine Learning explores the integration of quantum computing principles with classical learning algorithms to address complex computational challenges. Recent theoretical breakthroughs have demonstrated that QML models can possess a significant advantage in sample complexity compared to classical counterparts (huang2021power; huang2022quantum). Specifically, quantum models can extract effective information from high-dimensional feature spaces using significantly fewer training samples. This property is particularly relevant to our task of modeling individual creators, where the available ”ground truth” data is inherently scarce and high-dimensional, making classical deep learning prone to overfitting.
@@
While anomaly detection is a well-established field for identifying deviations from a learned norm (lavin2015evaluating3; garcia2009anomaly4; phua2010comprehensive5; schlegl2017unsupervised6), the application of QML to this domain remains nascent. Existing quantum approaches, such as kernel-based methods (belis2024quantum) or the Variational Quantum One-Class Classifier (VQOCC) (park2023variational), often struggle with hardware efficiency or lack experimental validation on general image datasets. To address these limitations, we adopt the PEQAD framework. Unlike prior methods, PEQAD employs an end-to-end variational circuit that maps multimodal features into a minimal hypersphere in the Hilbert space. Crucially, theoretical analysis and empirical results on superconducting processors suggest that PEQAD can achieve superior expressivity and detection performance with orders of magnitude fewer parameters than classical Deep SVDD (ruff2018deep). This extreme parameter efficiency makes it uniquely suited for detecting subtle semantic mutations in low-resource environments.
@@ 3 Methodology
Figure 2: Overview of the Quantum-BAR Framework. (1) Classical Fusion: Textual, audio, and visual features are extracted and fused into a vector $x$. (2) Quantum Encoder: The vector is amplitude-encoded into a 12-qubit system and processed by a VQC with parameters $\theta$. (3) Semantic Manifold: The VQC maps normal videos (blue dots) into a compact hypersphere centered at $|c\rangle$. Malicious mutations (red star) drift away from the center ($S_{quant}$) and align with the sensationalism axis ($S_{dir}$).
@@ 3.1 Multimodal Feature Extraction and Fusion
We represent each video $v$ by a set of modality-specific embeddings $\{f_{text}(v),f_{audio}(v),f_{visual}(v),f_{meta}(v)\}$. To construct a comprehensive semantic representation, we extract features across four distinct modalities. Textual features are derived from Automatic Speech Recognition (ASR) transcripts, which are processed using a pre-trained BERT encoder (gardazi2025bert) to capture the underlying semantic logic and linguistic patterns. Simultaneously, visual context is encoded through frame-level embeddings obtained via the CLIP model (radford2021learning), ensuring that visual fidelity is aligned with semantic content. Audio features incorporate prosodic elements, such as pitch and jitter, alongside spectral statistics; these are critical for identifying synthetic artifacts introduced by Text-to-Speech (TTS) systems, often used in malicious edits. Finally, we incorporate metadata features, including posting timestamps and tag usage distributions, to provide temporal and categorical context.
@@ 3.2.1 Amplitude Encoding
We utilize Amplitude Encoding to map the normalized classical vector $x$ into the amplitudes of a 12-qubit quantum state. This method allows us to encode $2^{n}$ classical features into $n$ qubits, providing exponential compression and capturing high-dimensional correlations that are computationally expensive for classical kernels[cite: 217, 130]. The encoded state is:
@@ 3.2.3 Hypersphere Optimization Objective
Following the Deep SVDD objective adapted for quantum circuits, we train the VQC to map all normal videos (the blogger’s authentic history) into a minimum-volume hypersphere in the Hilbert space, centered at a fixed state $|c\rangle$. The optimization objective is to minimize the expected distance between the mapped states and the center:
$\min_{\theta}\mathcal{L}_{PEQAD}=\frac{1}{N}\sum_{i=1}^{N}\big\||\psi_{out}(x_{i};\theta)\rangle-|c\rangle\big\|^{2}+\frac{\lambda}{2}\|\theta\|_{F}^{2},$ (4)
where $N$ is the batch size and $\lambda$ is a regularization term. By pulling all authentic semantic vectors toward $|c\rangle$, the VQC learns a compact manifold $\mathcal{M}$ where the creator’s logic and style are consistent.
@@ 3.3 Anomaly Scoring and Inference
Semantic mutations often drift not just away from the creator, but towards clickbait or conflict. We learn a global ”Sensationalism Direction” $\mathbf{w}_{mut}$ using LDA on a held-out corpus of known rumors and content-farm titles. The directional score is the cosine similarity:
@@
The final anomaly score combines the manifold distance (uncertainty) and the sensationalist drift (malicious intent):
@@ 4.1 Dataset Collection and Implementation
We manually curated a dataset sourced from two major short-video platforms, TikTok and Douyin. To ensure the validity of the manifold hypothesis, that a creator maintains a consistent stylistic and logical signature, we specifically selected 100 distinct bloggers (active from 2017–2024) with follower counts exceeding 100,000. These established accounts typically exhibit higher production stability and semantic consistency compared to smaller, variable accounts. The categories cover diverse topics, including news commentary, technology reviews, education, and entertainment. To simulate the real-world constraint of data scarcity, the training data for each blogger is strictly limited to approximately 20 authentic long-form videos (>2>2 minutes).

### v2

@@ 1 Introduction
The digital information ecosystem is currently witnessing a crisis of authenticity that transcends pixel-level forgery (Farid, 2022; Wardle and Derakhshan, 2017). While deepfake detection focuses on AI-synthesized artifacts, a more pervasive threat has emerged in the form of semantic mutation (Guarnera et al., 2020). In this phenomenon, unscrupulous actors, often operating as content farms, use traditional editing tools to reorder, decontextualize, or selectively splice genuine footage (Mubarak et al., 2023). A video subjected to these attacks remains visually authentic, containing the real speaker and genuine audio, but the underlying semantics and intent are fundamentally mutated to fabricate false narratives or clickbait (Kishwar et al., 2025; Luo et al., 2021).
The societal impact of such manipulation is profound and immediate  (Vaccari and Chadwick, 2020). A prominent example occurred on short-video platforms involving the top-tier technology creator MediaStorm (Yingshijufeng) [^1] For example, the widely discussed incident regarding unauthorized re-editing and platform content degradation reported in https://www.zhihu.com/question/1972852612236850378 (2025) or MediaStorm’s official response on Bilibili.. Malicious re-editing of their reviews resulted in a complete semantic inversion of the original critique, fabricating conflicts that did not exist. This incident not only dominated platform trending lists but also instigated widespread misinformation and targeted cyber-harassment against the creator. Such events underscore a critical vulnerability in the current media landscape: when visual fidelity is weaponized to distort factual truth, it can rapidly incite irrational public sentiment and cause irreversible reputational damage before the original context can be restored (Vosoughi et al., 2018; Vaccari and Chadwick, 2020). Detecting these high-level logic violations requires modeling the unique semantic manifold of the original creator (Ruff et al., 2018). A genuine blogger maintains a consistent latent signature defined by their linguistic patterns, logical structures, and prosodic delivery (Stamatatos, 2009). Ideally, an anomaly detector would learn the topology of this manifold and flag any video that drifts into low-density regions, indicating a deviation from the creator’s established persona.
@@
However, applying classical deep learning to this task faces a critical bottleneck: data scarcity (Huang et al., 2022, 2021; Wang et al., 2025a). A specific creator typically possesses a limited repository of high-quality, long-form original videos, often ranging from 20 to 50 samples. In our experiments, we further restrict the training set to approximately 20 videos per creator in order to simulate an extreme low-data regime, which reflects realistic deployment scenarios where only a small number of verified original videos are available. Classical anomaly detection methods, such as Deep Support Vector Data Description (Deep SVDD) (Ruff et al., 2018) or Variational Autoencoders (VAE) (Park et al., 2023; An and Cho, 2015), are notoriously data-hungry. When trained on such small, high-dimensional multimodal datasets, these models tend to suffer from severe overfitting or require massive parameter sets that fail to generalize to unseen semantic attacks (Wang et al., 2023). To surmount the obstacle of data scarcity, we turn to the emerging paradigm of Quantum Machine Learning (QML) (Biamonte et al., 2017). Recent theoretical breakthroughs have rigorously established that quantum models can extract physical information from complex, high-dimensional feature spaces more efficiently than their classical counterparts (Wang et al., 2025b). Notably, Huang et al. demonstrated that quantum learning algorithms can achieve a provable advantage in sample complexity, enabling accurate classification with significantly fewer training examples (Huang et al., 2021, 2022). This characteristic renders QML uniquely suitable for modeling the semantic manifolds of individual creators, where the “ground truth” is sparse yet highly complex.
Motivated by these findings, our approach integrates the Parameter-Efficient Quantum Anomaly Detection (PEQAD) algorithm into the domain of multimedia forensics (Wang et al., 2025a). Specifically, we fuse textual, audio, and visual features into a compact representation and employ a variational quantum circuit (VQC) to map normal blogger content into a minimal hypersphere in the quantum state space. This strategic integration yields a threefold enhancement. First, it mitigates overfitting by achieving robust anomaly detection using a few trainable quantum parameters, whereas comparable classical models often require hundreds. Second, the quantum state representation naturally models the entangled relationships between a blogger’s tone and logic (Gao et al., 2022). Third, the reduced model complexity facilitates potential deployment on near-term quantum devices or efficient simulators.
@@ 2 Related Work
Traditional VAD identifies abnormal events in surveillance feeds, typically employing autoencoders (Hasan et al., 2016) or multiple instance learning to detect deviations from normal motion patterns (Sultani et al., 2018). However, these methods rely heavily on visual cues (optical flow, gradients) (Liu et al., 2018) and are agnostic to linguistic semantics. They do not focus on detecting a video where a person is speaking calmly but saying something successfully spliced to reverse their original meaning (Abdalla et al., 2025).
@@
Research in misinformation has largely focused on “Fake News” classification via textual features or graph-based propagation patterns. Meanwhile, Deepfake detection analyzes frequency domain artefacts or blending boundaries in synthesized faces (Rössler et al., 2019). Our work addresses a distinct category: Cheapfakes or shallowfakes, where the manipulation is editorial rather than generative. As noted by (Paris and Donovan, 2019), this semantic manipulation is often more dangerous and harder to detect than AI-generated content. Crucially, existing deepfake detectors fail on these semantic alterations because no pixel-level artefacts are present.
@@
Recent works in NLP have explored logical consistency in text summarization (Maynez et al., 2020). We extend this notion to the multimodal domain, treating the blogger’s historical corpus as the “ground truth” reference for consistency checking.
Quantum Machine Learning.
Quantum Machine Learning explores the integration of quantum computing principles with classical learning algorithms to address complex computational challenges. Recent theoretical breakthroughs have demonstrated that QML models can possess a significant advantage in sample complexity compared to classical counterparts (Huang et al., 2021, 2022). Specifically, quantum models can extract effective information from high-dimensional feature spaces using significantly fewer training samples. This property is particularly relevant to our task of modeling individual creators, where the available ”ground truth” data is inherently scarce and high-dimensional, making classical deep learning prone to overfitting.
@@
While anomaly detection is a well-established field for identifying deviations from a learned norm (Lavin and Ahmad, 2015; Garcia-Teodoro et al., 2009; Phua et al., 2010; Schlegl et al., 2017), the application of QML to this domain remains nascent. Existing quantum approaches, such as kernel-based methods (Belis et al., 2024) or the Variational Quantum One-Class Classifier (VQOCC) (Park et al., 2023), often struggle with hardware efficiency or lack experimental validation on general image datasets. To address these limitations, we adopt the PEQAD framework. Unlike prior methods, PEQAD employs an end-to-end variational circuit that maps multimodal features into a minimal hypersphere in the Hilbert space. Crucially, theoretical analysis and empirical results on superconducting processors suggest that PEQAD can achieve superior expressivity and detection performance with orders of magnitude fewer parameters than classical Deep SVDD (Ruff et al., 2018). This extreme parameter efficiency makes it uniquely suited for detecting subtle semantic mutations in low-resource environments.
@@ 3 Methodology
Semantic representations of multimedia content often exhibit complex cross-modal correlations. For instance, the meaning of a statement may depend jointly on textual phrasing, vocal prosody, and visual context. Capturing such interactions using classical models typically requires high-dimensional feature expansions or deep neural networks with large numbers of parameters. Quantum feature maps provide an alternative mechanism for representing such correlations. By embedding classical feature vectors into a quantum Hilbert space through parametrized unitary transformations, variational quantum circuits implicitly generate highly nonlinear feature maps whose effective dimension grows exponentially with the number of qubits. Although the circuit itself may contain only a small number of trainable parameters, the resulting state representation can capture complex interactions among input features. In the context of semantic mutation detection, this property is particularly desirable. Subtle manipulations such as sentence reordering, contextual deletion, or mismatches between tone and content may correspond to high-order correlations between modalities rather than simple feature deviations. Quantum embeddings provide a compact mechanism to model these interactions while maintaining a small parameter footprint, making them suitable for creator-specific learning under limited data.
@@
Formally, let $\mathcal{M}_{c}\subset\mathbb{R}^{D}$ denote the semantic manifold associated with a specific creator $c$. Each authentic video $v_{i}$ is mapped to a multimodal feature vector $x_{i}\in\mathbb{R}^{D}$, and the set of historical content $\{x_{1},\dots,x_{N}\}$ forms a compact region approximating $\mathcal{M}_{c}$. Malicious edits introduce semantic perturbations that cause the embedding $\hat{x}$ of a manipulated video $\hat{v}$ to deviate from this region, producing a measurable anomaly signal.
@@
Figure 2: Overview of the Quantum-BAR Framework. (1) Classical Fusion: Textual, audio, and visual features are extracted and fused into a vector $x$. (2) Quantum Encoder: The vector is amplitude-encoded into a 12-qubit system and processed by a VQC with parameters $\theta$. (3) Semantic Manifold: The VQC maps normal videos (blue dots) into a compact hypersphere centered at $|c\rangle$. Malicious mutations drift away from the center ($S_{quant}$) and align with the sensationalism axis ($S_{dir}$).
@@ 3.1 Multimodal Feature Extraction and Fusion
We represent each video $v$ by a set of modality-specific embeddings $\{f_{text}(v),f_{audio}(v),f_{visual}(v),f_{meta}(v)\}$. To construct a comprehensive semantic representation, we extract features across four distinct modalities. Textual features are derived from Automatic Speech Recognition (ASR) transcripts, which are processed using a pre-trained BERT encoder (Gardazi et al., 2025) to capture the underlying semantic logic and linguistic patterns. Simultaneously, visual context is encoded through frame-level embeddings obtained via the CLIP model (Radford et al., 2021), ensuring that visual fidelity is aligned with semantic content. Audio features incorporate prosodic elements, such as pitch and jitter, alongside spectral statistics; these are critical for identifying synthetic artifacts introduced by Text-to-Speech (TTS) systems, often used in malicious edits. Finally, we incorporate metadata features, including posting timestamps and tag usage distributions, to provide temporal and categorical context.
@@ 3.2.1 Amplitude Encoding
We utilize Amplitude Encoding to map the normalized classical vector $x$ into the amplitudes of a 12-qubit quantum state. This method allows us to encode $2^{n}$ classical features into $n$ qubits, providing compression and capturing high-dimensional correlations that are computationally expensive for classical kernels (Wang et al., 2025a), although efficient state preparation remains an open challenge. The encoded state is:
@@ 3.2.3 Hypersphere Optimization Objective
Following the Deep SVDD objective adapted for quantum circuits, we train the VQC to map all normal videos (the blogger’s authentic history) into a minimum-volume hypersphere in the Hilbert space, centered at a fixed state $|c\rangle$. Following the PEQAD formulation (Wang et al., 2025a), the center state $|c\rangle$ is defined as the reference state representing the normal semantic region of the creator. In practice, $|c\rangle$ is determined according to the initialization strategy described in PEQAD, which ensures that the hypersphere center remains fixed during training while the circuit parameters $\theta$ adapt to minimize the distance between encoded samples and the center state. The optimization objective is to minimize the expected distance between the mapped states and the center:
$\min_{\theta}L_{PEQAD}=\frac{1}{N}\sum_{i=1}^{N}\left(1-|\langle c|\psi_{out}(x_{i};\theta)\rangle|^{2}\right)+\frac{\lambda}{2}\|\theta\|_{F}^{2}$ (4)
where $N$ is the batch size and $\lambda$ is a regularization term. This formulation corresponds to maximizing the quantum state fidelity between encoded samples and the center state. By pulling all authentic semantic vectors toward $|c\rangle$, the VQC learns a compact manifold $\mathcal{M}$ where the creator’s logic and style are consistent.
@@ 3.3 Anomaly Scoring and Inference
Semantic mutations often drift not just away from the creator, but towards clickbait or conflict. We learn a global “Sensationalism Direction” $w_{mut}$ using Linear Discriminant Analysis on a held-out corpus of clickbait headlines and rumor-related content collected from publicly available misinformation datasets. This direction captures semantic patterns commonly associated with sensationalized narratives. The directional score is the cosine similarity:
@@
The final anomaly score combines two complementary signals: structural deviation from the creator-specific semantic manifold and directional drift toward sensationalized content patterns. These two components capture different aspects of semantic mutation and are combined through a weighted linear score:
@@
3.4 Training and Inference Procedure
The complete Q-BAR pipeline is summarized in Algorithm 1. The procedure preserves the original BAR components, including multimodal fusion, classical manifold modelling, retrieval-based alignment, and directional mutation scoring, while augmenting them with the PEQAD quantum anomaly detector. In practice, the classical and quantum modules are trained separately and fused only at the score level, which improves stability in the low-data regime.
Algorithm 1 Q-BAR: Quantum-enhanced Blogger Anomaly Recognition
1:Input: creator-specific authentic dataset $\mathcal{D}=\{v_{i}\}_{i=1}^{N}$, query clip $\hat{v}$, external clickbait corpus $\mathcal{C}$
2:Multimodal feature extraction:
3:for each video $v_{i}\in\mathcal{D}$ do
4:  Extract textual, audio, visual, and metadata features
5:  Compute fused representation
6:  $f(v_{i})\leftarrow\mathrm{Fuse}\!\left(f_{\text{text}}(v_{i}),f_{\text{audio}}(v_{i}),f_{\text{visual}}(v_{i}),f_{\text{meta}}(v_{i})\right)$
7:end for
8:Classical manifold modelling:
9: Train a VAE on $\{f(v_{i})\}_{i=1}^{N}$ to obtain encoder $E$ and decoder $G$
10: Train a Deep SVDD model on $\{f(v_{i})\}_{i=1}^{N}$ to estimate a compact support region
11:Quantum manifold modelling:
12: Reduce fused features to a quantum-compatible dimension: $x_{i}\leftarrow r(f(v_{i}))$
[…]
14:for each training sample $x_{i}$ do
[…]
17:end for
18: Optimise $\theta$ by minimizing
$\mathcal{L}_{Q}=\frac{1}{N}\sum_{i=1}^{N}\left(1-\left|\langle c\mid\psi_{\mathrm{out}}(x_{i};\theta)\rangle\right|^{2}\right)+\frac{\lambda}{2}\|\theta\|_{F}^{2}$
19:Directional mutation modelling:
20: Learn the sensationalism direction $w_{\mathrm{mut}}$ from $\mathcal{C}$ using LDA
21:Inference for query clip $\hat{v}$:
22: Compute fused feature vector
[…]
24: Optionally retrieve candidate source videos using pHash and CLIP, then align them using DTW
25: Compute classical anomaly score
[…]
27: Compute quantum anomaly score
[…]
31: Compute directional mutation score
[…]
33: Normalize the three scores on the validation distribution:
[…]
35: Compute final anomaly score
[…]
36:if $S(\hat{v})>\tau$ then
37:  Flag $\hat{v}$ as a semantic anomaly
38:else
39:  Treat $\hat{v}$ as consistent with the creator manifold
40:end if
41:Output: final anomaly score S(v^)S($\hat${v})
Algorithm 1 highlights that PEQAD acts as an additional manifold learner in a quantum feature space, complementing the classical reconstruction-based detector. This design preserves the semantic interpretation of BAR while improving sensitivity to high-order cross-modal inconsistencies under limited training data.
@@ 4.1 Dataset Collection and Implementation
We manually curated a dataset sourced from two major short-video platforms, TikTok and Douyin. To ensure the validity of the manifold hypothesis, that a creator maintains a consistent stylistic and logical signature, we specifically selected 100 distinct bloggers (active from 2017–2024) with follower counts exceeding 100,000. These established accounts typically exhibit higher production stability and semantic consistency compared to smaller, variable accounts. The categories cover diverse topics, including news commentary, technology reviews, education, and entertainment. To simulate the real-world constraint of data scarcity, the training data for each blogger is strictly limited to approximately 20-50 authentic long-form videos (>2>2 minutes).
@@
Due to copyright and platform policy restrictions, the collected video dataset cannot be publicly released. However, the data collection protocol and preprocessing pipeline are described in detail to facilitate reproducibility.